OpenStudy (anonymous):

∫▒(sin⁡θ + sin⁡θ tan⁡〖^2 〗 )/sec⁡〖^2 θ〗 dθ intergrate ...help me please!

sam (.sam.):

Rewrite it properly I can't read it

OpenStudy (anonymous):

I=∫▒(sin⁡θ+sin⁡θ (tan⁡θ )^2)/(sec⁡θ )^2 dθ =∫▒[sec⁡θ tan⁡θ+(sin⁡θ )^3 ] dθ (sin⁡3θ=3 sin⁡θ-4(sin⁡θ )^3+c Or~4(sin⁡θ )^3=3 sin⁡θ-sin⁡3θ) =sec⁡〖θ-3 cos⁡θ 〗+cos⁡3θ/3+c

OpenStudy (anonymous):

correction 1/4 outside the bracket after sec theta.

OpenStudy (anonymous):

i have another question.. need your help..

OpenStudy (anonymous):

by the way, my answer for the integration question is -cos theta..... is it the correct answer...

OpenStudy (anonymous):

Let f(x)=sin2x f(x+h)=sin2(x+h)=sin(2x+2h)=sin2xcos2h+cos2xsin2h f(x+h)-f(x)=sin2xcos2h-sin2x+cos2xsin2h =sin2x(cos2h-1)+cos2xsin2h =sin2x(-2〖sin〗^2h)+cos2xsin2h =-2sin2x〖sin〗^2h+cos2x sin2h Divide by h (f(x+h)-f(x))/h=(-2sin2x〖sin〗^2 h)/h+cos2x sin2⁡h/2h*2 Taking limits as h approaches 0 lim┬(h=0)⁡〖(f(x+h)-f(x))/h=-2sin2x lim┬(h=0)⁡〖(〖sin〗^2 h)/h^2 〗 〗*h+2 cos⁡2x lim┬(h=0)⁡〖sin⁡2h/2h〗 d/dx {f(x)}=-2sin2x*1*0+2 cos⁡2x*1=2cos⁡2x d/dx sin⁡2x=2 cos⁡2x

OpenStudy (anonymous):

can you rewrite back the answer from *divide by h